OpenStudy (anonymous):

please show me the steps to solve this, A 1 1/2 oz. bottle of roll-on deodorant has a height of 4 in. and a radius of 3/4 inches. Find the surface area of the bottle if the bottle is shaped like a silo. Use 3.14 for π. Altitude of a launched object. The altitude of an object, in meters, is given by the polynomial h + vt − 4.9t2, where h is the height, in meters, at which the launch occurs, v is the initial upward speed (or velocity), in meters per second, and t is the number of seconds for which the object is airborne

OpenStudy (mertsj):

Well, a silo is shaped like a cylinder so I don't know why they couldn't have just said that. So it is composed of 2 circles and a rectangle. The two circles would be: \[2\pi (\frac{3}{4})^2\]

OpenStudy (mertsj):

The rectangle would have a base that is the circumference of the circle and it's height would be 4. So it's area would be: \[2\pi (\frac{3}{4})(4)\]

OpenStudy (mertsj):

Add the two together to get the total surface area.

OpenStudy (anonymous):

2*3.14(3/4) * (4) + 2*3.14*3/4^2 ?

OpenStudy (anonymous):

is that h ow it should be written?

OpenStudy (mertsj):

Just write the two parts like I did and put 3.14 for pi

OpenStudy (mertsj):

And put a plus sign between them.

OpenStudy (mertsj):

\[2(3.14)(\frac{3}{4})^2+2(3.14)(\frac{3}{4})(4)\]

OpenStudy (anonymous):

thank u